OpenStudy (anonymous):

Which of these pairs of points defines a line with a slope -1 over 3

OpenStudy (anonymous):

A) (6, -1) and (2, 5) B) (6, -1) and (5, 2) C) (-1, 6) and (5, 2) D) (-1, 6) and (2, 5)

OpenStudy (anonymous):

(5-6) = -1 2-(-1)=3 m= -1/3

OpenStudy (anonymous):

d

OpenStudy (anonymous):

D) (-1, 6) and (2, 5)